Every item in the Just a Dollar store is priced $1.00. When Jack opens the store, there is $150.00 in the cash register. When he counts the money in the cash register at the end of the day, the total is $719.75. If the sales tax rate is 7.5%, how many items were sold that day?
Answer by rfer(16322)

719.75-150=569.75
569.75/1.075=530
530/1=530 items